# Vision-and-Language Training Helps Deploy Taxonomic Knowledge but Does Not Fundamentally Alter It

### v1

@@ 5 H2: VLMs are better at deploying taxonomic knowledge
As mentioned in Section 4.1, solving a downstream task presupposes the domain knowledge recruited, and requires this knowledge to be correctly deployed in the context of the specific task. Hence, solving TaxonomiGQA requires (1) taxonomic knowledge and (2) its deployment specifically for scene description-based QA. Our analyses in the previous section did not show convincing evidence in support of the hypothesis that the underlying taxonomic knowledge differs substantially in our VLM-LM pairs. In light of this mostly negative result, we turn to our second hypothesis: VLMs are better at deployment of taxonomic knowledge. To test whether there is a difference when taxonomic knowledge is incorporated into the specific task context, we use contextual similarity of lexical representations and a Principal Component Analysis (PCA) of representations of questions. These analyses let us examine both the contextualized lexical representations as well as the holistic representation of the full question context. We use the Qwen2.5 pair in both analyses.
@@
To control for the confounding effect of the target label (Yes/No) when analyzing contextualized representations, we use a subset of TaxonomiGQA that has the same ground truth label for both positive and negative samples. In our dataset, this only includes cases where the ground truth answer is No. We further filter this dataset to instances where the models got the original, unsubstituted question right, and use the models’ Conditional Accuracy on substituted questions as the target of study. This gives us 37,790 and 40,145 individual samples for Qwen2.5-I and Qwen2.5-VL-I, respectively.
@@ 5.1 Contextualized representation similarity
Our first analysis aims to relate the behavioral outcome of a model for each question to the representational structure of the concepts in context. To this end, we investigate the contextualized representations of a target concept in the scene description in terms of their similarity to representations of its hypernym in the question (e.g., There is a doghypo on a yellow surfing board […]. In the scene, are there any mammalshyper?). The quantitative hypothesis is that greater contextualized hypernym-hyponym similarity (e.g., sim(dog, mammal) compared to hyponym similarity with negative samples (e.g., sim(dog, fruit) from There is a doghypo […]. In the scene, are there any fruitsneg?) would predict how well the model can answer the TaxonomiGQA questions. We use the 4 negative samples from TaxonomiGQA, and then fit a logistic regression model to predict model correctness (measured using the correct() function from Section 3.2) using the difference in cosine similarity between hypernym-hyponym, and maximum [^2] We note that taking the average instead of maximum results in substantially weaker trends. cosine similarity of 4 hyponym-negative sample pairs. Here, an odds ratio (of the difference term) being greater than 1 indicates that the similarity of the hyponym to hypernym (relative to negative samples) is more strongly associated with the model correctly answering questions, while the opposite is true if the odds ratio is less than 1. We perform this analysis using representations from every layer in the Qwen2.5 model pair, and take the maximum similarity in cases where the hyponym is mentioned in the scene more than once.
@@
Figure 3: Contextualized representational analysis on Qwen2.5-I and Qwen2.5-VL-I
Results
Figure 3(a) shows the layerwise odds ratios of the difference in similarities between concept pairs (sim-$\Delta$; discussed above), in predicting model correctness, for both models. For most layers, we see odds ratio values that are greater than 1.0, indicating positive association between sim-$\Delta$ and model correctness for both model classes. At the same time, the VLM odds ratios are often greater than those of the LM, with the LM odds ratios sometimes even veering off below the 1.0 level (which would suggest an association of sim-$\Delta$ with wrongness as opposed to correctness. Overall, this suggests that VL training helps establish stronger connections between model representations and behavior in task contexts requiring deployment of taxonomic knowledge.
@@ 5.2 Principal Component Analysis (PCA) of question representations
Like the previous analysis, we focus on the distinguishability of hypernym-hyponym relations from non-hypernym-hyponym relations, but consider whether this is captured in the holistic representation of the question context from data used in the previous section. Following Alhamoud et al. [1] (who tested negation sensitivity in VLMs), we take the last hidden state of the final layer of the text decoder to be the summary representation of the full context. Then we ask whether representations of questions that contain a hypernym-hyponym relation (e.g., dog-canine) are separated from representations of questions that contain a non-hypernym-hyponym relation (e.g., dog-bird) via PCA.
Figure 3(b) shows the first two principal components (PCs) of the question representations from the VLM & LM, with hypernym (green) vs. negative sample substitutions (orange) color coded. We see that the two types are largely visually distinct in both models, suggesting that their question representations do encode differences in terms of the taxonomic relations tested. To quantify (linear) separability, we fit a soft-margin support vector machine (SVM) classifier [7] on the first two principal components of the representations extracted from each model separately, and measure its error on the PC-representations—the greater the error, the poorer the separability. We find that the SVM error of the PCs of VLM representations is substantially lower than that of the LM, demonstrating that taxonomic distinctions are more linearly separable in VLM question representations. This complements the results from the previous analysis of contextualized embeddings, and suggests genuine differences in the representational states of the VLM and LM when the task contexts require taxonomic reasoning.
@@ 6 Why might vision training help?
[…]
@@ Results
[…]
@@ 7 Conclusion
Our statistical analyses do not provide causal evidence for the relation between behavior on TaxonomiGQA and the analyzed representations. Gaining causal evidence would require analyses more closely tied to the training data and objective, which is challenging due to the scale of the models as well as the scarcity of open data models. Nevertheless, we are excited to build on the dataset and methods presented here to conduct deeper analysis addressing the causal question in future work. Furthermore, our SVM-based separability analysis only sheds light on taxonomic distinctions that are linearly encoded, leaving room for future work to extend this to non-linear separability.
@@ Acknowledgments
We thank Yukyung Lee for her advice on creating better visualizations and Mahir Patel for earlier discussions on constructing the dataset, as well as for their general support. Pilot experiments for this work were conducted with the support of the PaliGemma Academic Program GCP Credit Award from Google awarded to the team. An in-person workshop partially dedicated to this work was funded by the Royal Netherlands Academy of Arts and Sciences (KNAW) Early Career Partnership. We acknowledge that the computational work reported on in this paper was performed on the Shared Computing Cluster which is administered by Boston University’s Research Computing Services, as well as on the burrata machine at TTIC.
@@ Appendix A Selected Model Pairs
Table 2 shows a list of model pairs used in this work, along with their metadata – parameters, modality, huggingface identifier, etc.
@@ Appendix B Extended Related Work
Appendix C Negative Sampling Details
@@
After dataset filtering, we identified 32 types of questions. 19 of them were excluded because substituting the object in the question with one not present in the scene could result in presupposition failures. For the remaining types, we sampled four negative objects for each question based on the following three criteria: the sampled argument is (1) not present in the scene, (2) not in the original argument’s hypernym chain, and (3) associated with the same set of attributes as the original arguments defined in GQA metadata. Due to inconsistencies and errors in the GQA metadata, we manually verified the attribute matches to ensure the naturalness and validity of each substitution. This process resulted in a final dataset consisting of 13 question types and reduced our taxonomy to 126 unique chains. Details of question types, examples, and statistics can be found in Table 3.
Appendix D VQA vs. Text and a Question-only Control
Table 4 shows results from evaluating VLMs on the original GQA questions across different formats: (1) the original VQA setup, conditioned on an image; (2) the Text setup, where they are conditioned on the scene description; and (3) a Question-only control where we condition them only on the question, without any context.
While it is difficult to compare between the VQA and the text setup, we see stark differences in the absolute values of the accuracies. The VLMs seem to answer the (positive sample, unsubstituted) questions with very high accuracy (sometimes near-perfect) relative to their performance on the subset of the VQA task we have used in this work. Next, the VLMs are substantially worse at the question-only baseline than they are in the text version, often times being closer to chance (50%). This question-only control is especially relevant for any potential concern readers might have about VQA data being present in the models’ training set. Since models are largely worse off at these relative to the text version of the dataset, the potential presence of VQA in the model’s training set is of little concern. One interesting observation here is that MLlama-3.2 (non-instruct tuned) performs similarly at the question only task and at the VQA task. This could suggest that it might not have been trained on VQA after all.
@@
Appendix E RSA Heatmaps
We depict heatmaps showing the pairwise cosine similarities computed for the transformed Unembedding vectors of the Qwen2.5 model pair, as well as pairwise path-similarity from WordNet, in Figure 5. The LM and VLM matrices look quite similar, while the WordNet matrices are more sparse, showing clearer depiction of hierarhical structure. We computed similar plots for all other models but left them out due to large file sizes. Full plots can be viewed on https://github.com/tinlaboratory/taxonomigqa.
@@ Appendix E RSA Heatmaps
Appendix F More Details about Visual Similarity Analysis
@@
Appendix G Experimental Resources
@@
Appendix H License Information

### v2

@@ 5 H2: VLMs are better at deploying taxonomic knowledge
As mentioned in Section 4.1, solving a downstream task presupposes the domain knowledge recruited, and requires this knowledge to be correctly deployed in the context of the specific task. Hence, solving TaxonomiGQA requires (1) taxonomic knowledge and (2) its deployment specifically for scene description-based QA. Our analyses in the previous section did not show convincing evidence in support of the hypothesis that the underlying taxonomic knowledge differs substantially in our VLM-LM pairs. In light of this mostly negative result, we turn to our second hypothesis: VLMs are better at deployment of taxonomic knowledge. To test whether there is a difference when taxonomic knowledge is incorporated into the specific task context, we used contextual similarity of lexical representations and a Principal Component Analysis (PCA) of representations of questions. These analyses let us examine both the contextualized lexical representations as well as the holistic representation of the full question context. We used the Qwen2.5 pair in both analyses.
@@
To control for the confounding effect of the target label (Yes/No) when analyzing contextualized representations, we used a subset of TaxonomiGQA that has the same ground truth label for both positive and negative samples. In our dataset, this only includes cases where the ground truth answer is No. We further filtered this dataset to instances where the models got the original, unsubstituted question right, and used the models’ Conditional Accuracy on substituted questions as the target of study. This gave us us 37,790 and 40,145 samples for Qwen2.5-I and Qwen2.5-VL-I, respectively.
@@ 5.1 Contextualized representation similarity
Our first analysis aims to relate the behavioral outcome of a model for each question to the representational structure of the concepts in context. To this end, we investigated the contextualized representations of a target concept in the scene description in terms of their similarity to representations of its hypernym in the question (e.g., There is a doghypo on a yellow surfing board […]. In the scene, are there any mammalshyper?). The quantitative hypothesis is that greater contextualized hypernym-hyponym similarity (e.g., sim(dog, mammal) compared to hyponym similarity with negative samples (e.g., sim(dog, fruit) from There is a doghypo […]. In the scene, are there any fruitsneg?) would predict how well the model can answer the TaxonomiGQA questions. We used the 4 negative samples from TaxonomiGQA, and then fit a logistic regression model to predict model correctness (measured using the correct() function from Section˜3.2) using the difference in cosine similarity between hypernym-hyponym, and maximum [^2] We note that taking the average instead of maximum results in substantially weaker trends. cosine similarity of 4 hyponym-negative sample pairs. Here, an odds ratio (of the difference term) being greater than 1 indicates that the similarity of the hyponym to hypernym (relative to negative samples) is more strongly associated with the model correctly answering questions, while the opposite is true if the odds ratio is less than 1. We performed this analysis using representations from every layer in the Qwen2.5 model pair, and took the maximum similarity in cases where the hyponym is mentioned in the scene more than once.
@@
Figure 3: Contextualized representational analysis on Qwen2.5-I and Qwen2.5-VL-I.
Figure˜3(a) shows the layerwise odds ratios of the difference in similarities between concept pairs (sim-$\Delta$; discussed above), in predicting model correctness, for both models. For most layers, we see odds ratios greater than 1.0, indicating a positive association between sim-$\Delta$ and model correctness for both model classes. At the same time, the VLM odds ratios are often greater than those of the LM, with the LM odds ratios sometimes even veering off below the 1.0 level (which would suggest an association of sim-$\Delta$ with wrongness as opposed to correctness. Overall, this suggests that VL training helps establish stronger connections between model representations and behavior in task contexts requiring deployment of taxonomic knowledge.
@@ 5.2 Principal Component Analysis (PCA) of question representations
Like in the previous analysis, we focused on the distinguishability of hypernym-hyponym relations from non-hypernym-hyponym relations, but considered whether this is captured in the representation of the question context from data used in the previous section. Following Alhamoud et al. [1] (who tested negation sensitivity in VLMs), we took the last hidden state of the final layer of the text decoder to be the summary representation of the full context. Then we asked whether representations of questions that contain a hypernym-hyponym relation (e.g., dog-canine) are separated from representations of questions that contain a non-hypernym-hyponym relation (e.g., dog-bird) via PCA.
Figure˜3(b) shows the first two principal components (PCs) of the question representations from the VLM & LM, with hypernym (green) vs. negative sample substitutions (orange) color coded. We see that the two types are largely visually distinct in both models, suggesting that their question representations do encode differences in terms of the taxonomic relations tested. To quantify (linear) separability, we fit a soft-margin support vector machine (SVM) classifier [7] on the first two principal components of the representations extracted from each model separately, and measured its error on the PC-representations—greater the error, the poorer the separability. We found that the SVM error of the PCs of VLM representations is substantially lower than that of the LM, demonstrating that taxonomic distinctions are more linearly separable in VLM question representations. This complements the results from the previous analysis of contextualized embeddings, and suggests genuine differences in the representational states of the VLM and LM when the task contexts require taxonomic reasoning.
5.3 On the distinction between knowledge possession and deployment
Collectively, our results highlight that VL training does not change the underlying taxonomic knowledge within LMs, but rather affects its deployment in task contexts that require sensitivity to taxonomic knowledge. We see two specific reasons why this distinction could be important.
First, storing or representing knowledge differs from learning its “functional consequences” [46]. A model may robustly encode category information (e.g., that robins are birds), yet fail to recruit this knowledge when the context demands it. TaxonomiGQA is designed precisely to probe this aspect of deployment: in order to be successful, a model must not only store taxonomic knowledge, but use it appropriately when answering questions. Practically, teasing apart knowledge possession and deployment can inform decisions about (post-)training data selection: if a model’s limitations stem from representational gaps, additional encyclopedic knowledge may help; if the issue lies in deployment, more diverse contextual supervision, involving contexts in which such knowledge is recruited, could be more effective. This distinction can also motivate solutions; for instance transferring task vectors from models that are better at deployment [16].
Second, the distinction has implications for cognitive and philosophical interpretations of multimodal learning—in particular, for drawing appropriate conclusions about the roles of linguistic versus (added) extralinguistic exposure. For instance, the platonic representation hypothesis [20] suggests that models trained on sufficiently large amounts of data converge toward similar internal representations, irrespective of modality. Our findings provide a complementary perspective to this hypothesis. While independent unimodal models might converge to similar taxonomic representations, combining information from multiple modalities can result in non-trivial changes that go beyond representational convergence (in our case, in terms of how knowledge is accessed and deployed).
@@ 6 Why might vision training help?
To test this hypothesis, we first estimated hypernym-hyponym visual similarities by computing the cosine similarity between the image representation of a leaf node object and the image representations of other objects within its parent node (i.e., its hypernym) for concepts in our taxonomy. The image representations are extracted from the target VLM’s (Qwen2.5-VL-I) vision encoder. Importantly, the images themselves are sourced from an independent dataset (THINGS [15]) so that our conclusion is not tied to specific images in GQA. Rather, they are intended as estimates of visual similarity more broadly. More details about the image similarity computation is in Appendix˜H. Then, we tested the extent to which this similarity predicts Conditional Accuracy of the VLM on hypernym questions where it outperforms its LM counterpart, using a linear mixed-effects model. Specifically, we predicted Conditional Accuracy of the VLM between each hyponym-hypernym pair using the pair’s visual similarity as a fixed effect, and included random slopes and intercepts for each hypernym (model formula: $\texttt{cond\_acc}\sim\texttt{viz\_sim}+(1+\texttt{viz\_sim}\mid\texttt{hypernym})$).
@@ Results
We found a significant global effect of visual similarity in predicting Conditional Accuracy ($b=0.52,\textrm{SE}=0.19,p<.01$). These results are much weaker when using the text-only LM’s Conditional Accuracy as the dependent variable ($b=0.23,\textrm{SE}=0.17,p=0.18$), suggesting that image similarity captures interesting properties related to the success of the VLM and uniquely so for VLMs. We also found interesting hypernym-specific random effects, where the effect of similarity varies greatly depending on the hypernym. Figure˜4 shows the random slopes for each hypernym. This substantial individual variation aligns with our initial intuition that visual similarity would help some relations more than others. To quantify this intuition concretely, we annotated the higher level concepts in our taxonomy in terms of the % of the time the visual similarity to their hyponyms were above the median. This is meant to capture the difference between equine and animal we discussed earlier: members of equine are more similar to each other, more so than members of animal are (i.e., visually cohesive). The colors of the bars in Figure˜4 are mapped to the degree of visual cohesion, where darker bars mean more cohesive. We see that the degree of cohesion generally lines up with effect sizes of similarity on predicting VLM performance, with mostly lighter bars on the left edge and darker bars on the right edge. The figure zooms into the concepts on either edge (left: animal, right: band), showing a sample of images corresponding to those concepts to illustrate the low visual cohesion of animal and high visual cohesion of band. Overall, the results present a promising lead into elucidating the source of improvement in VLMs, establishing a potential link between visual similarity, visual cohesion, and behavioral QA performance.
@@ 7 Conclusion
Our analyses do not provide causal evidence for the relation between behavior on TaxonomiGQA and the analyzed representations. Gaining causal evidence would require analyses more closely tied to the training data and objective, which is challenging due to the scale of the models as well as the scarcity of open data models. Additionally, a caveat to our results is that VL-tuned models do encounter more text-data in addition to visual supervision. This confound can be teased apart in future work by training LMs on the text-only portion of the VL training data. Furthermore, our SVM-based separability analysis is only applicable to taxonomic distinctions that are linearly encoded, leaving room for future work to extend this to non-linear separability.
@@ Acknowledgments
We thank Yukyung Lee for her advice on creating better visualizations and Mahir Patel for earlier discussions on constructing the dataset, as well as for their general support. We also thank the anonymous NeurIPS reviewers and the Area Chair for helpful feedback. Pilot experiments for this work were conducted with the support of the PaliGemma Academic Program GCP Credit Award from Google awarded to the team. An in-person workshop partially dedicated to this work was funded by the Royal Netherlands Academy of Arts and Sciences (KNAW) Early Career Partnership. We acknowledge that the computational work reported on in this paper was performed on the Shared Computing Cluster which is administered by Boston University’s Research Computing Services, as well as on the burrata machine at TTIC. Kanishka Misra is supported by the Donald D. Harrington Faculty Fellowship at UT Austin.
@@ Appendix A Selected Model Pairs
Table˜2 shows a list of model pairs used in this work, along with their metadata – parameters, modality, huggingface identifier, etc.
@@ Appendix B Extended Related Work
Appendix C Taxonomy Filtering and Annotation
Before manual annotation and removal of specific chains, we first identified 52 highly abstract concepts [^3] These concepts are listed in the GQA Hierarchies First Pass.ipynb notebook located under notebooks/ in our repository. (e.g. entity, conveyance, act, etc.) to be removed from all chains. After generating the initial hypernym chains, we conducted a second round of manual inspection to identify and address cases of “non-ideal” categorizations, defined as instances where either (a) the assigned hypernym was not the canonical category of the object (e.g., bubble categorized as circle), or (b) the chain consisted solely of abstract elements that were missed during the first filtering step. Through this process, we identified 611 problematic cases. Of these, 296 were corrected by querying WordNet for alternative hypernym chains, leaving 315 unfixable cases; these were subsequently removed (as reported in the main text).
Appendix D Negative Sampling Details
@@
After dataset filtering, we identified 32 types of questions. 19 of them were excluded because substituting the object in the question with one not present in the scene could result in presupposition failures. For the remaining types, we sampled four negative objects for each question based on the following three criteria: the sampled argument is (1) not present in the scene, (2) not in the original argument’s hypernym chain, and (3) associated with the same set of attributes as the original arguments defined in GQA metadata. Due to inconsistencies and errors in the GQA metadata, we manually [^4] We remove attributes that introduce non-standard property attribution, such as ”happy trees”, ”swimming flowers”, ”fluffy apple”. verified the attribute matches to ensure the naturalness and validity of each substitution. This process resulted in a final dataset consisting of 13 question types and reduced our taxonomy to 126 unique chains. Details of question types, examples, and statistics can be found in Table˜3.
Appendix E VQA vs. Text and a Question-only Control
Table˜4 shows results from evaluating VLMs on the original GQA questions across different formats: (1) the original VQA setup, conditioned on an image; (2) the Text setup, where they are conditioned on the scene description; and (3) a Question-only control where we condition them only on the question, without any context.
While it is difficult to compare between the VQA and the text setup, we see stark differences in the absolute values of the accuracies. The VLMs seem to answer the (positive sample, unsubstituted) questions with very high accuracy (sometimes near-perfect) relative to their performance on the subset of the VQA task we have used in this work. Next, the VLMs are substantially worse at the question-only baseline than they are in the text version, often times being closer to chance (50%). This question-only control is especially relevant for any potential concern readers might have about VQA data being present in the models’ training set. Since models are largely worse off at these relative to the text version of the dataset, the potential presence of VQA in the model’s training set is of little concern. One interesting observation here is that MLlama-3.2 (non-instruct tuned) performs similarly at the Question-only task and at the VQA task. This could suggest that it might not have been trained on VQA after all.
@@
Appendix F Supplementary Experiment on the Rodriguez et al. Dataset
Property inheritance testing datasets such as those introduced by Rodriguez et al. [56] involve attributing a novel, nonsense property (e.g., is daxable) to concepts given that their parents are attributed with it—e.g., Given that birds are daxable, is it true that robins are daxable? Answer with Yes or No. This particular dataset includes a robustness check in the form of a single negative sample per concept and includes 2016 positive samples, spanning 44 superordinate categories and 1281 subordinate categories. [^5] This dataset has a larger set of categories than TaxonomiGQA, as its taxonomy is not constrained by visual scenes. We evaluated our model pairs on this dataset with minimal prompt tuning, [^6] We added the quantifier “all” to the premise - e.g., Given that all birds are daxable, is it true that robins are daxable? Answer with Yes or No, to make the logical inference more coherent. This modification improved the QA accuracy significantly, particularly for Qwen2.5-VL-I and MLlama-3.2. and found that, as shown in Table 5, all VLMs (except MLlama-3.2 and Llava-OV) outperformed their LM counterparts, reinforcing the robustness of the observed VLM >> LM trend on models’ behavioral performances on taxonomic understanding within a QA setting.
Table 5: Performance comparison between LMs and VLMs on the Rodriguez et al. dataset. $\Delta(\text{VLM--LM})$ denotes the difference in accuracy between the VLM and its corresponding LM.
Pair LM VLM $\bm{\Delta(\text{VLM--LM})}$
Llama-3.1 vs. MLlama-3.2 0.50 0.68 0.18
Llama-3.1-I vs. MLlama-3.2-I 0.51 0.50 -$0.01$
Mistral-v0.2-I vs. Llava-Next 0.69 0.78 0.09
Qwen2 vs. Molmo-D 0.60 0.81 0.21
Qwen2-I vs. Llava-OV 0.84 0.81 -$0.02$
Qwen2.5-I vs. Qwen2.5-VL-I 0.78 0.83 0.05
Vicuna vs. Llava-1.5 0.50 0.74 0.24
Appendix G RSA Heatmaps
We depict heatmaps showing the pairwise cosine similarities computed for the transformed Unembedding vectors of the Qwen2.5 model pair, as well as pairwise path-similarity from WordNet, in Figure˜5. The LM and VLM matrices look quite similar, while the WordNet matrices are more sparse, showing clearer depiction of hierarhical structure. We computed similar plots for all other models but left them out due to large file sizes. Full plots can be viewed on https://github.com/tinlaboratory/taxonomigqa.
@@ Appendix G RSA Heatmaps
Appendix H More Details about Visual Similarity Analysis
@@
Appendix I Experimental Resources
@@
Appendix J License Information